A box has dimensions that have consecutive numbers for its length, width and height, the height having the smallest of the three dimensions. If the length and width are increased by 2 cm and the height is doubled, the volume now becomes 192 cm3 greater than the original box. What are the dimensions of the original box?
Answer by stanbon(75887)

A box has dimensions that have consecutive numbers for its length, width and height, the height having the smallest of the three dimensions. If the length and width are increased by 2 cm and the height is doubled, the volume now becomes 192 cm3 greater than the original box. What are the dimensions of the original box?
--------------
Let the dimensions be "x-1", x, and "x+1"
-------------------
Original dimensions:
height = x-1
length = x
width = x+1
------------------
Equation:
Original volume is (x-1)*x*(x+1) = x(x^2-1) = x^3-x
--------------------
New dimensions:
height = 2x-2
length = x+2
width = x+3
-----------------
Equation:
new volumn - old volumn = 19 2cm^3
2(x-1)(x+2)(x+3) - (x^3-x) = 192
x = 4
----------
Answer:
height = x-1 = 3
length = x = 4
width = x+1 = 5
=======================
Cheers,
Stan H.